What force must be exerted on the master cylinder of a hydraulic lift to support the weight of a 2500-kg car (a large car) resting on the slave cylinder? The master cylinder has a 4.00-cm diameter and the slave has a 25.0-cm diameter.

1
Expert's answer
2022-02-09T16:15:27-0500

Given:

"m=2500\:\rm kg"

"d_1=\rm 4.00\: cm"

"d_2=\rm 25.0\: cm"


The hydraulic press equation says

"\frac{F_1}{A_1}=\frac{F_2}{A_2}"

Hence, the required force

"F=mg\frac{A_1}{A_2}=2500*9.8*\frac{(4.00)^2}{(25.0)^2}=627\: \rm N"